### v1

@@ 2 Methodology
$\boldsymbol{\alpha}$ $=(\alpha_{1},...,\alpha_{p})$
$\boldsymbol{\gamma}$ $=(\gamma_{1},...,\gamma_{p})$
$\boldsymbol{\psi}$ $=(\psi_{1},...,\psi_{L})$
$\boldsymbol{Z}$ $=(z_{1,1},...,z_{1,n_{1}},z_{2,1},...,z_{2,n_{2}},...,z_{p,n_{p}})$
$\boldsymbol{\beta}$ $=\begin{pmatrix}\beta_{11}&\ldots&\beta_{1L}\\ \vdots&\ddots&\vdots\\ \beta_{p1}&\ldots&\beta_{pL}=\end{pmatrix}=\begin{pmatrix}\beta_{1}\\ \vdots\\ \beta_{p}\end{pmatrix}$
where the full conditional for $\gamma_{p}$ for all $p$ in 1,…,$P$ is distributed as $\text{Gamma}(\gamma_{0}k+n_{p},k+|I_{p}|)$, for $\psi_{l}$ with $l=1,..,L$ we have
$\pi(\psi_{l}|\cdot)\propto\prod_{p=1}^{P}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big)f_{\text{NIW}}(\psi_{l})$
which by conjugated analysis is distributed as $\text{NIW}(\mu_{l},\eta_{l},\Sigma_{l},\nu_{l})$ where
$\mu_{l}$ $=\frac{\eta\mu_{0}+m_{l}\bar{y}}{\eta+m_{l}}$
$\eta_{l}$ $=\eta+m_{l}$
$\nu_{n}$ $=\nu+m_{l}$
$\Sigma_{l}$ $=\Sigma_{0}+S+\frac{\eta m_{l}}{\eta+m_{l}}(\bar{y}-\mu_{0})(\bar{y}-\mu_{0})$
@@
$m_{l}$ $=\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}\mathbbm{1}(z_{p,i}=l)$
$\bar{y}$ $=\frac{\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}(x_{p,i},y_{p,i})\mathbbm{1}(z_{p,i}=l)}{m_{l}}$
$S$ $=\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{t}}((x_{p,i},y_{p,i})-\bar{y})((x_{p,i},y_{p,i})-\bar{y})^{T}\mathbbm{1}(z_{p,i}=l),$
@@
$\pi(\alpha_{1}|\cdot)$ $\propto\frac{e^{-\alpha_{1}}\alpha_{1}^{\alpha_{0}}\alpha_{2}^{\alpha_{1}}}{\Gamma(\alpha_{1})}\frac{\displaystyle\prod_{l=1}^{L}\beta_{1l}^{\frac{\alpha_{1}}{L}}}{\text{B}(\frac{\alpha_{1}}{L},\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L})}$
$\pi(\alpha_{P}|\cdot)$ $\propto e^{-\alpha_{P}}\alpha_{P}^{\alpha_{P-1}}\frac{\displaystyle\prod_{l=1}^{L}\beta_{Pl}^{\alpha_{P}\beta_{P-1,l}}}{\text{B}(\alpha_{P}\beta_{P-1,1},...,\alpha_{P}\beta_{P-1,L})}$
@@
$\pi(\alpha_{p}|\cdot)$ $\propto\frac{e^{-\alpha_{p}}\alpha_{p}^{\alpha_{p-1}}\alpha_{p+1}^{\alpha_{p}}}{\Gamma(\alpha_{p})}\frac{\displaystyle\prod_{l=1}^{L}\beta_{pl}^{\alpha_{p}\beta_{p-1,l}}}{\text{B}(\alpha_{p}\beta_{p-1,1},...,\alpha_{p}\beta_{p-1,L})}$
@@
$\pi(\beta_{1}|\cdot)\propto\prod_{l=1}^{L}\beta_{1l}^{m_{1l}+\frac{\alpha_{1}}{L}-1}\frac{\displaystyle\prod_{l=1}^{L}\beta_{2l}^{\alpha_{2}\beta_{1,l}}}{\text{B}(\alpha_{2}\beta_{1,1},...,\alpha_{2}\beta_{1,L})}$
meanwhile for $p$ in 2,,P2,...,P we have $\beta_{p}|\cdot\sim\text{Beta}(m_{pl}+\alpha_{p}\beta_{pl},...,m_{pl}+\alpha_{p}\beta_{pl})$ where $m_{pl}=\displaystyle\sum_{i=1}^{n_{p}}\mathbbm{1}(z_{p,i}=l)$, finally
$\mathbb{P}(Z_{pi}=l|\cdot)\propto\beta_{pl}\phi_{\psi_{l}}(x_{p,i},y_{p,i})$
[…]
As discussed in [chakrabarti2024graphical], the total conditional distributions of $\alpha_{i}$, i=1,,Pi=1,...,P, and $\beta_{1}$ are not standard distributions and are the main bottleneck in the MCMC. Since all $\alpha_{i}$ are positive random variables, we proposed a random walk movement to explore them. According to [chakrabarti2024graphical], we used proposals based on the SALTSampler introduced by [director2017efficient] to explore $\beta_{1}$. Then, we opted for a Hybrid MCMC [robert1999monte] to sample the posterior distribution; the code is available at github.com/isaiasmanuel/NHPP.
@@
$\chi_{ii}$ $=\begin{cases}1,&\text{if earthquake $i$ is a background event}\\ 0,&\text{otherwise}\end{cases}$ (8)
$\chi_{ij}$ $=\begin{cases}1,&\text{if earthquake $i$ is an aftershock of $j$}\\ 0,&\text{otherwise}\end{cases},$ (9)
where s=0,,ms=0,...,m. These variables represent the branch structure of the earthquakes, and the log-likelihood of the ETAS model incorporating them is given by
@@
where $I^{2}$ denotes $\{1,2,...,N\}\times\{1,2,...,N\}$, being $\times$ the cartesian product, $\ell_{\text{O}}(\cdot)$ is the log-likelihood due to the offspring, and $\ell_{\text{I}}(\cdot)$ is the log-likelihood due to immigrants. They are given by
@@
However, in our case, this evaluation is not necessary to estimate the background intensity function because the features of the [sawires2019updated] catalog allow us to simply assume that all the $\chi_{ij}$ are 0 for all $i\neq j$, whereas $\chi_{ii}=1$ for all $i$.
@@ 3.1 Simulated Data
$\lambda(x,y,t)=\Big(50\mathbbm{1}(t<5)+100\mathbbm{1}(t\geq 5)\Big)\Big(h(t)g_{1}(x,y)+(1-h(t))g_{2}(x,y)\Big)$ (12)
@@
$h(t)$ $=\frac{1}{1+e^{-\frac{t-T}{2}}}$
$g_{1}(x,y)$ $=\frac{2}{3}\phi_{(0,0),\mathbf{1}_{2\times 2}}(x,y)+\frac{1}{3}\phi_{(2,2),\mathbf{1}_{2\times 2}}(x,y)$
$g_{2}(x,y)$ $=\frac{2}{3}\phi_{(6,2),\mathbf{1}_{2\times 2}}(x,y)+\frac{1}{3}\phi_{(4,6),\mathbf{1}_{2\times 2}}(x,y)$
The idea behind the proposed $\lambda$ function is to create a spatio-temporal density function that can capture the typical challenges observed in an earthquake catalog. For example, the number of recorded earthquakes may increase due to changes in instrumentation or aseismic transients as we discussed in Section 1.
In this example we take for the Normal-inverse-Wishart distribution $\mu_{0}=(1,1)$, $\Sigma_{0}=\mathbb{I}_{2}$, $\nu=3$ and $\eta=0.1$, where $\mathbb{I}_{2}$ is the $2\times 2$ identity matrix, the idea is to have a non informative prior distribution for the parameters in the mixture. For all the $\gamma_{p}$ we take $\gamma_{0}=70$ and $k=0.1$, which results in a vague prior distribution and $\alpha_{0}=1$. Also, a regular partition of the interval $(0,T)$ was used with $P=8$ and $L=8$.
@@
Figure 2 shows the intensity function in the whole space at different times. This visual representation facilitates the observation of spatial and temporal changes. The realization of this NHPP was generated using the thinning algorithm [lewis1979simulation2, illian2008statistical]. The simulated catalog and the code to generate it can be consulted in github.com/isaiasmanuel/NHPP.
@@
Figure 3 presents the posterior mean of $\gamma_{p}f_{p}$, as defined in Equation (5). Then, to utilize the information in the posterior distribution more effectively, we present the Figure 4. The colors of the figure are the same as in Figure 3, but now we include a transparency based on the posterior coefficient of variation (CV), which is defined by
@@
As can be seen in Figure 4, we have lower coefficient of variation values in regions with higher values of $\lambda$ (i.e., and as expected, the regions with the highest number of points are better estimated). To complement Figure 4, we also present Figure 5, where the histograms of the marginal posterior samples are presented for each $\gamma_{p}$ and the real value $\gamma_{p}$ (vertical dashed blue line), according to Equation (12). Figure 5 reveals that the posterior distributions of $\gamma_{p}$ adequately contain the real value for all the $p=1,2,$....
Figure 5: Histogram of the posterior $\gamma_{p}$, p=1,2,,Pp=1,2,...,P using simulated data. Blue dashed lines are the real $\gamma_{p}$ for each interval.
Figure 6: Clustering of the simulated data using 4 groups.
In this work, we do not focus on the inference related to each element in the mixture. Instead, we are interested in inferring the intensity function $\lambda$. Nevertheless, it is worth mentioning that we observed the label switching problem, as expected in a finite mixture problem [green1995reversible, zhang2004learning, bouguila2012fully]. For this reason, we have implemented the algorithm proposed by [papastamoulis2010artificial] (the code is available at github.com/isaiasmanuel/NHPP), mainly to select the thinning in the chain.
Finally, to validate our model in Figure 6, we present the clustering of the observed points. We show the mean of the Gaussian distributions (red crosses) used to define $\lambda$ in Equation 12. As can be seen, we recover the four mixture components adequately, even when the mixture weights change over time. The clusters were taken using the least-square criterion proposed in [dahl2006model], as our affinity matrix for spectral clustering [von2007tutorial], and the number of clusters was taken using the eigengap heuristic.
From this example, we have seen how GDP can be used to accurately estimate the intensity function of an inhomogeneous spatial Poisson process. That is why, in the next section, we apply this same methodology to the [sawires2019updated] catalog to recover the background seismic function.

### v2

@@ 2 Methodology
$\boldsymbol{\alpha}$ $=(\alpha_{1},...,\alpha_{P}),$
$\boldsymbol{\gamma}$ $=(\gamma_{1},...,\gamma_{P}),$
$\boldsymbol{\psi}$ $=(\psi_{1},...,\psi_{L}),$
$\boldsymbol{Z}$ $=(z_{1,1},...,z_{1,n_{1}},z_{2,1},...,z_{2,n_{2}},...,z_{P,n_{P}}),$
$\boldsymbol{\beta}$ $=\begin{pmatrix},\beta_{11}&\ldots&\beta_{1L}\\ \vdots&\ddots&\vdots\\ \beta_{P1}&\ldots&\beta_{PL}=\end{pmatrix}=\begin{pmatrix}\beta_{1}\\ \vdots\\ \beta_{P}\end{pmatrix}$,
To facilitate efficient sampling from the posterior distribution, we derive the full conditional distributions of the model parameters using the definition of conditional probability. In particular, the full conditional distribution of $\gamma_{1}$ is given by
$f(\gamma_{1}|\alpha,\boldsymbol{\beta},\boldsymbol{\gamma}_{-1},\boldsymbol{\psi},\mathbf{Z},X)=\frac{f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z}|X)}{f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}_{-1},\boldsymbol{\psi},\mathbf{Z},X)}\propto f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z}|X),$
where $\boldsymbol{\gamma}_{-1}$ denotes the vector $\boldsymbol{\gamma}$ without its first component, i.e., the full conditional distribution of a parameter is proportional to the posterior distribution, treating the remaining parameters as constants. Therefore, it is easy to see that the full conditional for $\gamma_{p}$ for all p=1,,Pp=1,...,P is distributed as $\text{Gamma}(\gamma_{0}k+n_{p},k+|I_{p}|)$.
For $\psi_{l}$ with $l=1,..,L$ the full conditional is given by
$\pi(\psi_{l}|\cdot)\propto\prod_{p=1}^{P}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big)f_{\text{NIW}}(\psi_{l}),$
which corresponds to independent and identically distributed (i.i.d.) Gaussian samples, which is a famous result of conjugated analysis that can be consulted in [murphy2022probabilistic], and is distributed as $\text{NIW}(\mu_{l},\eta_{l},\Sigma_{l},\nu_{l})$ where
$\mu_{l}$ $=\frac{\eta\mu_{0}+m_{l}\widehat{\mu}_{xy}}{\eta+m_{l}},$
$\eta_{l}$ $=\eta+m_{l},$
$\nu_{n}$ $=\nu+m_{l},$
$\Sigma_{l}$ $=\Sigma_{0}+S+\frac{\eta m_{l}}{\eta+m_{l}}(\widehat{\mu}_{xy}-\mu_{0})(\widehat{\mu}_{xy}-\mu_{0})^{T},$
@@
$m_{l}$ $=\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}\mathbbm{1}(z_{p,i}=l),$
$\widehat{\mu}_{xy}$ $=\frac{\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}(x_{p,i},y_{p,i})\mathbbm{1}(z_{p,i}=l)}{m_{l}},$
$S$ $=\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}((x_{p,i},y_{p,i})-\widehat{\mu}_{xy})((x_{p,i},y_{p,i})-\widehat{\mu}_{xy})^{T}\mathbbm{1}(z_{p,i}=l).$
It is important to note that the $\psi_{l}$ for l=1,,Ll=1,...,L are not independent as it can be seen from Equation (8) even when they were assumed a priori to be independent. Furthermore we want to highlight that their full conditional means take into account the samples $(x_{p,i},y_{p,i})$ for all $i=1,...,n_{p}$ and p=1,2,,Pp=1,2,...,P, i.e. they are including information from the observed earthquakes among all the partitions.
@@
$\pi(\alpha_{1}|\cdot)$ $\propto\frac{e^{-\alpha_{1}}\alpha_{1}^{\alpha_{0}-1}\alpha_{2}^{\alpha_{1}}}{\Gamma(\alpha_{1})}\frac{\displaystyle\prod_{l=1}^{L}\beta_{1l}^{\frac{\alpha_{1}}{L}}}{\text{B}(\frac{\alpha_{1}}{L},\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L})},$
$\pi(\alpha_{P}|\cdot)$ $\propto e^{-\alpha_{P}}\alpha_{P}^{\alpha_{P-1}-1}\frac{\displaystyle\prod_{l=1}^{L}\beta_{Pl}^{\alpha_{P}\beta_{P-1,l}}}{\text{B}(\alpha_{P}\beta_{P-1,1},...,\alpha_{P}\beta_{P-1,L})},$
@@
$\pi(\alpha_{p}|\cdot)$ $\propto\frac{e^{-\alpha_{p}}\alpha_{p}^{\alpha_{p-1}-1}\alpha_{p+1}^{\alpha_{p}}}{\Gamma(\alpha_{p})}\frac{\displaystyle\prod_{l=1}^{L}\beta_{pl}^{\alpha_{p}\beta_{p-1,l}}}{\text{B}(\alpha_{p}\beta_{p-1,1},...,\alpha_{p}\beta_{p-1,L})}.$
@@
$\pi(\beta_{1}|\cdot)\propto\prod_{l=1}^{L}\beta_{1l}^{m_{1l}+\frac{\alpha_{1}}{L}-1}\frac{\displaystyle\prod_{l=1}^{L}\beta_{2l}^{\alpha_{2}\beta_{1,l}}}{\text{B}(\alpha_{2}\beta_{1,1},...,\alpha_{2}\beta_{1,L})},$
in a similar way, for p=1,2,,P1p=1,2,...,P-1
$\pi(\beta_{p}|\cdot)\propto\prod_{l=1}^{L}\beta_{pl}^{m_{pl}+\alpha_{p}\beta_{p-1,l}-1}\frac{\displaystyle\prod_{l=1}^{L}\beta_{p+1,l}^{\alpha_{p+1}\beta_{p,l}}}{\text{B}(\alpha_{p+1}\beta_{p,1},...,\alpha_{p+1}\beta_{p,L})},$
where $m_{pl}=\displaystyle\sum_{i=1}^{n_{p}}\mathbbm{1}(z_{p,i}=l)$. Meanwhile for $p=P$, since the term $p+1$ does not exist, we have
$\beta_{P}|\cdot\sim\text{Dirichlet}(m_{P1}+\alpha_{P}\beta_{P-1,1},...,m_{Pl}+\alpha_{P}\beta_{P-1,L}).$
It is important to highlight that the full conditional distribution of $\beta_{p}$, for $p=2,\dots,P-1$, depends on the neighboring vectors $\beta_{p-1}$ and $\beta_{p+1}$. Thus, even though an autoregressive structure is specified a priori, the resulting posterior dependence reinforces temporal coupling between adjacent coefficients. In combination with the values of $\boldsymbol{\alpha}$, this structure allows the model to capture smooth temporal behavior when $\boldsymbol{\alpha}$ is large, while permitting more abrupt changes in the density when $\boldsymbol{\alpha}$ is close to zero.
A limitation of this approach is that, as the norm of the time partition tends to zero, the number of observations within each interval becomes very small, which complicates estimation. However, this issue can be mitigated by imposing an informative prior on $\boldsymbol{\alpha}$, thereby enforcing a smoothness structure over the spatio-temporal evolution. This regularization may be particularly useful in seismic catalogs where large-magnitude slow slip events (SSEs) are observed without the occurrence of strong earthquakes.
Finally,
$\mathbb{P}(Z_{pi}=l|\cdot)\propto\beta_{pl}\phi_{\psi_{l}}(x_{p,i},y_{p,i}),$
for $i=1,2,...,n_{p}$.
The autoregressive GDP of order 1 was selected for simplicity, however, as detailed in [chakrabarti2024graphical] this framework can be extended to greater orders. Furthermore, as can be seen from the posterior distribution, the parameters are not independent even for non-contiguous time instants.
As discussed in [chakrabarti2024graphical], the full conditional distributions of $\alpha_{i}$, i=1,,Pi=1,...,P, and $\beta_{p}$, for p=1,,P1p=1,...,P-1 are not standard distributions and are the main bottleneck in the MCMC. Since all $\alpha_{i}$ are positive random variables, we propose a random walk step to explore them. According to [chakrabarti2024graphical], we used proposals based on the SALTSampler introduced by [director2017efficient] to explore $\beta_{p}$. Then, we opted for a Hybrid MCMC [robert1999monte] to sample the posterior distribution; the code is available at github.com/isaiasmanuel/NHPP.
@@
$\chi_{ii}$ $=\begin{cases}1,&\text{if earthquake $i$ is a background event}\\ 0,&\text{otherwise}\end{cases},$ (9)
$\chi_{ij}$ $=\begin{cases}1,&\text{if earthquake $i$ is an aftershock of $j$}\\ 0,&\text{otherwise}\end{cases}.$ (10)
These variables represent the branch structure of the earthquakes, and the log-likelihood of the ETAS model incorporating them is given by
@@
where $I^{2}$ denotes $\{1,2,...,N\}\times\{1,2,...,N\}$, where $\times$ denotes the Cartesian product, $\ell_{\text{O}}(\cdot)$ is the log-likelihood due to the offspring, and $\ell_{\text{I}}(\cdot)$ is the log-likelihood due to immigrants. They are given by
@@
However, in our case, this evaluation is not necessary to estimate the background intensity function because the features of the [sawires2019updated] catalog allow us to assume simply that all the $\chi_{ij}$ are 0 for all $i\neq j$, whereas $\chi_{ii}=1$ for all $i$.
@@ 3.1 Simulated Data
$\lambda(x,y,t)=\Big(50\mathbbm{1}(t<5)+100\mathbbm{1}(t\geq 5)\Big)\Big(h(t)g_{1}(x,y)+(1-h(t))g_{2}(x,y)\Big),$ (13)
@@
$h(t)$ $=\frac{1}{1+e^{-\frac{t-T}{2}}},$
$g_{1}(x,y)$ $=\frac{2}{3}\phi_{(0,0),\mathbf{1}_{2\times 2}}(x,y)+\frac{1}{3}\phi_{(2,2),\mathbf{1}_{2\times 2}}(x,y),$
$g_{2}(x,y)$ $=\frac{2}{3}\phi_{(6,2),\mathbf{1}_{2\times 2}}(x,y)+\frac{1}{3}\phi_{(4,6),\mathbf{1}_{2\times 2}}(x,y).$
The idea behind the proposed $\lambda$ function is to construct a spatio-temporal intensity function capable of capturing the typical challenges observed in an earthquake catalog. For example, in [banales2025inhomogeneous], it has been shown that SSEs can modify seismicity. In particular, in southern Mexico, SSEs can last longer than one year, which motivates the function $h$ in Equation (13), allowing for a smooth transition between modes. Additionally, the number of recorded earthquakes may increase due to changes in instrumentation or aseismic transients, as discussed in Section 1. In Mexico, improvements in instrumentation, as described in [vergnolle2010slow, cruz2018seismogeodetic], have increased the number of recorded earthquakes. This non-smooth change is reflected in Equation (13) by changing the productivity parameter from 50 to 100 at time 5.
In this example we take for the Normal-inverse-Wishart distribution $\mu_{0}=(1,1)$, where the vector (1,1) was chosen arbitrarily to be inside $D$, $\Sigma_{0}=\mathbb{I}_{2}$, $\nu=3$ and $\eta=0.1$, where $\mathbb{I}_{2}$ is the $2\times 2$ identity matrix, the idea is to have a non informative prior distribution for the parameters in the mixture. For all the $\gamma_{p}$ we take $\gamma_{0}=70$ and $k=0.1$, which results in a vague prior distribution and $\alpha_{0}=1$. Also, a regular partition of the interval $(0,T)$ was used with $P=8$ and $L=8$.
@@
Figure 2 shows the intensity function in the whole space at different times. This visual representation facilitates the observation of spatial and temporal changes. The stochastic simulation of this NHPP was generated using the thinning algorithm [lewis1979simulation2, illian2008statistical], which is summarized as follows
  1. 1.
    Define $\lambda^{*}$ as the maximum of $\lambda(\cdot),$ i.e.
    $\lambda^{*}:=\max_{(x,y)\in D,t\in(0,T)}\lambda(x,y,t).$
  2. 2.
    Simulate a sample $\{(x_{i}^{s},y_{i}^{s},t_{i}^{s})\}_{i=1}^{N^{*}}$ that follows a homogeneous Poisson Process (HPP) in $D$ over $(0,T)$ with intensity $\lambda^{*}$. For a deeper discussion about how to simulate an HPP consult [bremaud2020point].
  3. 3.
    For each $i$ in $\{1,2,...,N^{*}\}$, the sample $(x_{i}^{s},y_{i}^{s},t_{i}^{s})$ is accepted as a valid simulation of the NHPP with intensity function $\lambda(\cdot)$ with probability
    $p_{i}=\frac{\lambda(x_{i}^{s},y_{i}^{s},t_{i}^{s})}{\lambda^{*}}.$
The simulated catalog, consisting of 727 events, and the code to generate it can be consulted in github.com/isaiasmanuel/NHPP.
@@
Figure 3 presents the posterior mean of $\gamma_{p}f_{p}$, as defined in Equation (5). Then, to utilize the information in the posterior distribution more effectively, we present Figure 4. The colors of the figure are the same as in Figure 3, but now we include a transparency based on the posterior coefficient of variation (CV), which is defined by
@@
As can be seen in Figure 4, we have lower coefficient of variation values in regions with higher values of $\lambda$ (i.e., as expected, the regions with the highest number of points are better estimated). To complement Figure 4, we also present Figure 5, where the histograms of the marginal posterior samples are presented for each $\gamma_{p}$ and the true value $\gamma_{p}$ (vertical dashed blue line), according to Equation (13). Figure 5 reveals that the posterior distributions of $\gamma_{p}$ adequately contain the true value for all p=1,2,,Pp=1,2,...,P.
Figure 5: Histogram of the posterior $\gamma_{p}$, p=1,2,,Pp=1,2,...,P using simulated data, blue dashed lines are the real $\gamma_{p}$ for each interval
Figure 6: 4-groups clustering of the simulated data
In this work, we do not focus on the inference related to each element in the mixture. Instead, we are interested in inferring the intensity function $\lambda$. Nevertheless, it is worth mentioning that we observed the label switching problem, as expected in a finite mixture problem [green1995reversible, zhang2004learning, bouguila2012fully]. For this reason, we have implemented the algorithm proposed by [papastamoulis2010artificial] (the code is available at github.com/isaiasmanuel/NHPP), and we also relabel by lexicographic order of the cluster mean in order to select the thinning in the chain. To read the discussion of the burn-in and thinning of the chain, please see the supplementary material.
To validate our model, we present in Figure 6 the clustering of the observed points. We show the mean of the Gaussian distributions (red crosses) used to define $\lambda$ in Equation (13). This clustering was obtained using the variation of information metric (VI) introduced by [meilua2007comparing] and the library mcclust.ext presented in [wade2018bayesian]. As can be seen, we recover the four mixture components adequately, even when the mixture weights change over time.
We also performed clustering using Binder’s loss, as proposed in [binder1978bayesian, dahl2006model]. Nevertheless, we observed a substantial overestimation of the number of clusters, as it can be consoulted in the supplementary material, which is consistent with the discussions in [wade2018bayesian, dahl2022search]. For this reason, we prefer to use VI. Finally, we would like to mention the papers [socher2011spectral, duan2024spectral], where new priors for clusters are proposed based on distance measures and Bayesian spanning forests that rely on spectral clustering, which may be of interest to readers primarily focused on clustering analysis.
From this example, we have seen how GDP can be used to accurately estimate the intensity function of an inhomogeneous spatial Poisson process. That is why, in the next section, we apply this same methodology to the [sawires2019updated] catalog to recover the background seismicity function.